Assume that an investor buys 100 shares of stock at $48 per share, putting up a 58% margin. a. What is the debit balance in this transaction? b. How much equity capital must the investor provide to make this margin transaction? a. The debit balance in this transaction is . (Round to the nearest dollar.)

Answer A and B Please
Answer #1

Answer A - Debit Balance in the transaction is $2016

Reason -

Securities price = $48 * 100 share = $4800

Debit balance = securities price ( 1 - margin )

Debit balance = $4800 ( 1 - 0.58 )

Debit Balance = $4800 * 0.42 = $2016

Answer B - Margin = $2784

Reason -

Margin = securities price * margin rate

Margin = $4800 * 0.58 = $2784.

If the solution helped, please give it a thumbs up. Thank you.